Falconmaster:

Two numbers have prime factorizations of 2 2 • 3 • 5 and 2 • 3 2 • 5. Which expressions will give the greatest common factor and the least common multiple? GCF 2 • 3; LCM 2 • 3 • 5 GCF 2^2 • 3^2; LCM 2^2 • 3^2 • 5 2 GCF 2 • 3 • 5; LCM 2 2 • 3^2 • 5 GCF 2^2 • 3^2 • 5; LCM 2 • 3

Zarkon:

I assume the two numbers are \(2^2\cdot 3\cdot 5\) and \(2\cdot 3^2\cdot 5\)

Falconmaster:

yes

Zarkon:

the GCD (GCF as you call it) ar just the numbers they have in common...in this case one 2 one 3 and one 5 ie \(2\cdot 3\cdot 5\) the LCM are the numbers with the highest powers between the two numbers ie \(2^2\cdot 3^2\cdot 5\)